# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ I-A Background
The increasing prevalence of Internet of things (IoT) devices and the demand for real-time applications have propelled a shift from centralized cloud computing to a distributed edge-cloud paradigm [1, 2, 3]. While cloud computing initially offers scalability and cost-effectiveness, limitations such as latency and bandwidth constraints spur the emergence of edge computing, bringing computation closer to data sources [4, 5, 6]. Consequently, edge-cloud collaborative computing (ECCC) arises, integrating the strengths of both paradigms to facilitate efficient resource utilization and enable complex AI tasks across the edge-cloud continuum [7, 8]. AI integration within ECCC systems addresses the inherent complexity of distributed environments, enabling optimized resource allocation and strategic model deployment for enhanced system performance [9]. In addition, AI empowers intelligent decision-making at the edge, enabling real-time responses. However, realizing distributed intelligence in ECCC systems introduces challenges, including efficient model partitioning and resource-aware model optimization, along with ensuring seamless edge-cloud collaboration and addressing security and privacy concerns [10, 11].
@@
Technical foundations underlying ECCC span multiple critical domains that require comprehensive understanding. Model optimization represents a fundamental challenge in deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency and performance [12]. Model compression techniques including pruning, quantization, and knowledge distillation, alongside adaptive learning approaches such as transfer learning, federated learning, and continual learning enable models to evolve with changing data distributions [6]. Resource management in distributed edge-cloud environments involves complex orchestration of computational, communication, and storage resources across heterogeneous infrastructure, requiring intelligent task offloading strategies, dynamic resource allocation algorithms, and energy-efficient optimization techniques to balance performance, cost, and sustainability [13]. Security and privacy protection mechanisms are paramount in ECCC systems due to the distributed nature of data processing and the involvement of multiple stakeholders, necessitating advanced cryptographic protocols, privacy-preserving machine learning techniques, and secure multi-party computation approaches to protect sensitive information while enabling collaborative intelligence [14].
@@
ECCC has emerged as a crucial paradigm to support the growing demands of real-time, intelligent applications, particularly with the proliferation of artificial intelligence of things (AIoT) devices [15]. However, inherent limitations in edge devices, such as constrained computational power and limited battery life, hinder their ability to independently handle computationally intensive AI workloads [16]. Additionally, the dynamic nature of workloads in these environments requires sophisticated resource allocation and task offloading strategies to maintain performance and prevent bottlenecks [17]. Applications like autonomous driving [18] and industrial automation [19] further necessitate low latency and high reliability, demanding intelligent resource management and optimization [20]. Consequently, AI offers a powerful set of tools, including deep reinforcement learning (DRL) for dynamic resource allocation [21] and multi-objective optimization for task scheduling [22], to address these challenges and improve the efficiency and performance of ECCC systems. For instance, federated learning enables collaborative model training across distributed devices while preserving privacy [16], ultimately contributing to enhanced quality of service (QoS) by optimizing resource utilization and minimizing latency.
@@
Communication network evolution plays a pivotal role in enabling effective edge-cloud collaboration by providing the foundational infrastructure that connects distributed computational resources. Recent advancements in network technologies, particularly software-defined networking (SDN) [23] and network function virtualization (NFV) [24], have significantly enhanced the flexibility and programmability of communication infrastructures supporting ECCC. Advanced technologies facilitate dynamic network resource allocation and traffic engineering, which are essential for optimizing data flows between edge devices and cloud servers [25]. Furthermore, emerging wireless communication paradigms, including millimeter-wave communications, massive MIMO, and non-orthogonal multiple access (NOMA) [26], are revolutionizing the capacity, reliability, and efficiency of edge-cloud connectivity. The integration of network slicing techniques [27] enables the creation of virtualized, isolated network segments with tailored quality-of-service guarantees for diverse edge-cloud applications. Moreover, cross-layer optimization approaches that jointly consider communication, computation, and caching resources [28] are becoming increasingly important for minimizing end-to-end latency and maximizing system throughput. As communication networks evolve toward 6G, the synergistic optimization of networking and computing resources will be instrumental in addressing the growing demands of distributed intelligence across the edge-cloud continuum [29].
@@
The selection and optimization of communication protocols significantly impact the performance, security, and reliability of edge-cloud collaborative systems [30]. Protocol optimization encompasses multiple layers with distinct considerations: (1) Application-layer protocols such as HTTP/HTTPS remain fundamental for web-based edge services, while MQTT and CoAP provide lightweight alternatives specifically designed for resource-constrained IoT devices with intermittent connectivity [31]. MQTT’s publish-subscribe model offers excellent scalability for sensor data aggregation scenarios, whereas CoAP’s REST-based approach provides seamless integration with existing web infrastructures, though MQTT introduces broker dependency that can create single points of failure, while CoAP’s UDP-based nature may require additional mechanisms for reliable delivery in lossy networks [32]; (2) Network-layer protocols including IPv6 and emerging protocols like named data networking (NDN) address the scalability challenges of massive IoT deployments. IPv6 provides abundant address space essential for large-scale edge device connectivity, while NDN’s content-centric approach enables efficient data distribution and caching across edge networks, nevertheless IPv6 adoption remains inconsistent across heterogeneous edge environments, and NDN requires significant infrastructure changes that may limit near-term deployment [33]; and (3) Transport-layer innovations such as QUIC and specialized protocols like BBR congestion control offer improved performance over traditional TCP in high-latency, variable-bandwidth edge-cloud connections [34]. QUIC’s built-in encryption and reduced connection establishment overhead particularly benefit mobile edge scenarios, while BBR’s model-based congestion control adapts better to the diverse network conditions encountered in edge deployments. The choice of appropriate protocols depends on specific application requirements, device capabilities, and network characteristics, with optimal edge-cloud systems often employing protocol stacks that dynamically adapt to changing conditions and requirements [35].
@@ I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, each with distinct scopes and technical emphases. Early works like Abbas et al. [36] and Ferrer et al. [37] establish a foundational understanding of edge computing and decentralized cloud architectures respectively, with their primary focus on architectural foundations rather than AI-driven model optimization. Similarly, Liu et al. [38] provide valuable insights into edge computing systems and tools, while Wang et al. [39] contribute significantly to offloading algorithms, with their emphasis primarily on different aspects of edge-cloud integration rather than comprehensive AI technique integration. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [40] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management, with their primary emphasis on system-level considerations rather than model optimization techniques. Domain-specific surveys like Arthurs et al. [41] focus on intelligent transportation systems with specialized depth in this domain. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain an active area of exploration, with Qu et al. [42] providing a dedicated investigation into mobile edge intelligence for LLMs, focusing primarily on resource-efficient techniques within their specialized scope. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications with their emphasis on domain-specific considerations.
@@
The most relevant contemporary works include Yang et al. [4], who extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, with their work emphasizing network dynamics as their primary contribution. Similarly, the comprehensive works by Yao et al. [11], Gu et al. [43], and Wang et al. [2] each make valuable contributions with distinct emphases. Specifically, [11] provides excellent coverage of edge-cloud collaboration and AI techniques with particular strength in these core areas. [43] delivers strong analysis of AI-enhanced cloud-edge-terminal networks with good application coverage, emphasizing network architecture aspects. [2] offers comprehensive model optimization coverage across the end-edge-cloud spectrum with particular depth in optimization techniques.
@@
Two surveys are particularly relevant to our work and warrant detailed comparison. Yang et al. [4] provides a comprehensive survey of reinforcement learning approaches for mobile edge computing with excellent coverage of resource management and autonomous driving applications. Their work has a specific focus that differs from our survey in several key aspects: (1) Specialized computing paradigm emphasis, while they focus extensively on edge computing and resource management, our work provides comprehensive coverage of cloud integration and edge-cloud collaboration frameworks; (2) Domain-specific technical focus, their emphasis on reinforcement learning, while thorough, represents a specialized approach compared to our comprehensive treatment of model optimization techniques such as compression, adaptation, and neural architecture search that are central to modern ECCC systems; (3) Different scope for privacy and security, their work emphasizes other technical aspects, while our survey provides systematic attention to privacy-preserving mechanisms and security enhancement for real-world ECCC deployments; (4) Specialized application coverage, while strong in autonomous driving, our work provides comprehensive coverage of manufacturing and healthcare applications that increasingly rely on edge-cloud collaboration. Qu et al. [42] presents a focused investigation into mobile edge intelligence for large language models with strong coverage of LLMs, privacy considerations, and healthcare applications. Their work complements our survey with different emphases: (1) Specialized technical scope, while excellent in LLM-specific optimizations, our survey provides systematic coverage of general model optimization techniques applicable across diverse AI models and architectures; (2) Mobile edge intelligence focus, their focus on mobile edge intelligence represents a specialized area, while our work emphasizes comprehensive edge-cloud collaboration frameworks and distributed intelligence paradigms; (3) Specialized application domain, while strong in healthcare applications, our survey provides detailed analysis of manufacturing and industrial IoT scenarios that represent major application domains for ECCC; (4) Domain-specific security analysis, their privacy coverage addresses specific requirements, while our work provides comprehensive treatment of security mechanisms and adversarial considerations essential for robust ECCC systems.
@@
[36] 2018 Mobile Edge Comp., Network Architecture
[37] 2019 Decentralized Cloud, Ad Hoc Computing
[38] 2019 Edge Computing, Systems, Tools
[39] 2019 Edge Cloud, Offloading Algorithms
[40] 2021 Edge Computing, ECC, CPS
[41] 2022 EC Computing, ITS, Connected Vehicles
[3] 2022 Edge Intel., URLLC, Digital Twin
[44] 2023 Offloading, Federated Cloud-Edge-Fog
[45] 2023 Edge Computing, Sensor-Cloud
[11] 2023 EC Polarization, Collab. AI, Edge Comp.
[46] 2024 AIGC, Mobile Networks, EC Generative AI
[43] 2024 AI, CET Network, Collaborative Computing
[2] 2024 DL, Distributed Computing, Model Opt.
[4] 2025 RL, Mobile Edge Comp., Res. Mgmt.
[42] 2025 Mobile Edge Intel., LLMs, Res. Mgmt.
@@ I-C Scope and Contributions
Figure 1: Overall structure of this survey. The framework is organized into five layers: (1) Foundations & architecture covering introduction and edge-cloud fundamentals, (2) technical implementation focusing on model optimization and resource management techniques, (3) protection framework addressing privacy and security enhancement, (4) practical deployment discussing applications and performance analysis, and (5) research vision outlining challenges and future directions.
As illustrated in Fig. 1, we provide a systematic and multi-layered examination of ECCC, addressing the critical intersection of distributed intelligence and model optimization. Edge-cloud systems face unprecedented challenges due to heterogeneous infrastructures, resource limitations, and growing computational demands from AI applications, making a comprehensive review of this rapidly evolving field both timely and necessary.
Our survey establishes a comprehensive taxonomy that organizes ECCC into five interconnected layers, each addressing specific aspects of distributed intelligence deployment. Foundations layer encompasses architectural paradigms, distributed intelligence frameworks, and fundamental resource constraints that govern system design. Technical implementation layer focuses on core enabling technologies including model optimization techniques (compression, adaptation, neural architecture search) and AI-driven resource management strategies (task offloading, resource allocation, energy optimization). Protection framework layer addresses critical security and privacy concerns through advanced cryptographic protocols, privacy-preserving learning, and secure computation techniques. Practical deployment layer examines real-world applications across diverse domains and comprehensive performance analysis methodologies. Finally, the Research vision layer identifies current challenges and future research directions that will shape the evolution of ECCC systems. Our taxonomic framework provides researchers and practitioners with a structured understanding of how different components interact within the broader ECCC ecosystem, enabling systematic development and optimization of distributed intelligent systems.
By systematically analyzing ECCC architectures, resource management strategies, and performance metrics, the survey bridges existing gaps in understanding how AI can enhance decentralized computing environments. Moreover, it investigates critical enabling technologies such as model compression, adaptation, resource-aware optimization, and security mechanisms, which are pivotal for overcoming current limitations. The scope spans from architectural foundations and optimization techniques to practical applications and future research directions, offering researchers and practitioners a holistic perspective of this transformative paradigm.
@@
The three primary technical categories of ECCC each address distinct challenges and exhibit specific applicability constraints. Model optimization techniques (Sec. III) primarily target the fundamental challenge of deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency. Model optimization approaches are particularly suitable for scenarios involving deep learning model deployment, privacy-sensitive distributed training, and real-time inference applications. However, their effectiveness is constrained by the inherent trade-offs between model accuracy and compression ratios, the heterogeneity of edge device capabilities, and the complexity of managing distributed learning processes. Model compression methods excel in reducing computational overhead but may compromise accuracy, while federated learning approaches provide privacy preservation at the cost of increased communication complexity and convergence challenges with non-IID data distributions. AI-driven resource management strategies (Sec. IV) focus on optimizing resource utilization and task orchestration across the heterogeneous edge-cloud continuum. Resource management approaches are most effective in scenarios with dynamic workloads, multi-tenant environments, and applications requiring intelligent load balancing across diverse computational resources. The primary limitations include the computational complexity of multi-objective optimization problems, the difficulty in predicting system behavior under varying conditions, and scalability challenges in large-scale deployments. While task offloading mechanisms can significantly reduce latency and energy consumption, they require sophisticated algorithms to handle dependency relationships and may introduce communication overhead that negates performance benefits. Privacy and security enhancement frameworks (Sec. V) address the critical requirements for protecting sensitive data and ensuring system trustworthiness in distributed environments. Security techniques are essential for healthcare, financial, and government applications where regulatory compliance and data protection are paramount. Nevertheless, privacy-preserving mechanisms often introduce computational overhead and complexity in key management, while security measures must continuously evolve to address emerging threats in heterogeneous edge environments.
@@
  • First Comprehensive Survey. To the best of our knowledge, this is the first comprehensive survey that specifically targets the intersection of distributed intelligence and model optimization within broader field of AI-driven edge-cloud collaborative computing, reviewing recent advances across architectures, techniques, and security.

### v5

@@ I-A Background
The increasing prevalence of Internet of things (IoT) devices and the demand for real-time applications have propelled a shift from centralized cloud computing to a distributed edge-cloud paradigm [masaracchia20256genabled, wang2024endedgecloud, vanhuynh2022edge]. While cloud computing initially offers scalability and cost-effectiveness, limitations such as latency and bandwidth constraints spur the emergence of edge computing, bringing computation closer to data sources [yang2025edge, moreschini2024edge, yang2025growthadaptive]. Consequently, edge-cloud collaborative computing (ECCC) arises, integrating the strengths of both paradigms to facilitate efficient resource utilization and enable complex AI tasks across the edge-cloud continuum [khalyeyev2023characterization, sun2025latencyaware]. AI integration within ECCC systems addresses the inherent complexity of distributed environments, enabling optimized resource allocation and strategic model deployment for enhanced system performance [wang2025efficient]. In addition, AI empowers intelligent decision-making at the edge, enabling real-time responses. However, realizing distributed intelligence in ECCC systems introduces challenges, including efficient model partitioning and resource-aware model optimization, along with ensuring seamless edge-cloud collaboration and addressing security and privacy concerns [liu2023distributed, yao2023edgecloud].
@@
Technical foundations underlying ECCC span multiple critical domains that require comprehensive understanding. Model optimization represents a fundamental challenge in deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency and performance [sharma2024deep]. Model compression techniques including pruning, quantization, and knowledge distillation, alongside adaptive learning approaches such as transfer learning, federated learning, and continual learning enable models to evolve with changing data distributions [yang2025growthadaptive]. Resource management in distributed edge-cloud environments involves complex orchestration of computational, communication, and storage resources across heterogeneous infrastructure, requiring intelligent task offloading strategies, dynamic resource allocation algorithms, and energy-efficient optimization techniques to balance performance, cost, and sustainability [lu2024a2cdrl]. Security and privacy protection mechanisms are paramount in ECCC systems due to the distributed nature of data processing and the involvement of multiple stakeholders, necessitating advanced cryptographic protocols, privacy-preserving machine learning techniques, and secure multi-party computation approaches to protect sensitive information while enabling collaborative intelligence [seifelnasr2024privacypreserving].
@@
ECCC has emerged as a crucial paradigm to support the growing demands of real-time, intelligent applications, particularly with the proliferation of artificial intelligence of things (AIoT) devices [tuli2023ai]. However, inherent limitations in edge devices, such as constrained computational power and limited battery life, hinder their ability to independently handle computationally intensive AI workloads [zhou2024enhancing]. Additionally, the dynamic nature of workloads in these environments requires sophisticated resource allocation and task offloading strategies to maintain performance and prevent bottlenecks [paul2025quantumenhanced]. Applications like autonomous driving [poularakis2019joint] and industrial automation [ren2021synergy] further necessitate low latency and high reliability, demanding intelligent resource management and optimization [jansen2023specrg]. Consequently, AI offers a powerful set of tools, including deep reinforcement learning (DRL) for dynamic resource allocation [xu2024fusion] and multi-objective optimization for task scheduling [zhang2024manyobjective], to address these challenges and improve the efficiency and performance of ECCC systems. For instance, federated learning enables collaborative model training across distributed devices while preserving privacy [zhou2024enhancing], ultimately contributing to enhanced quality of service (QoS) by optimizing resource utilization and minimizing latency.
@@
Communication network evolution plays a pivotal role in enabling effective edge-cloud collaboration by providing the foundational infrastructure that connects distributed computational resources. Recent advancements in network technologies, particularly software-defined networking (SDN) [guo2024madrlom] and network function virtualization (NFV) [bensalah2024vnf], have significantly enhanced the flexibility and programmability of communication infrastructures supporting ECCC. Advanced technologies facilitate dynamic network resource allocation and traffic engineering, which are essential for optimizing data flows between edge devices and cloud servers [bensalem2023scaling]. Furthermore, emerging wireless communication paradigms, including millimeter-wave communications, massive MIMO, and non-orthogonal multiple access (NOMA) [letaief2022edge], are revolutionizing the capacity, reliability, and efficiency of edge-cloud connectivity. Moreover, the integration of network slicing techniques [li2021slicingbased] enables the creation of virtualized, isolated network segments with tailored quality-of-service guarantees for diverse edge-cloud applications. Moreover, cross-layer optimization approaches that jointly consider communication, computation, and caching resources [kanduri2024edgecentric] are becoming increasingly important for minimizing end-to-end latency and maximizing system throughput. As communication networks evolve toward 6G, the synergistic optimization of networking and computing resources will be instrumental in addressing the growing demands of distributed intelligence across the edge-cloud continuum [kokkonen2023autonomy].
@@
The selection and optimization of communication protocols significantly impact the performance, security, and reliability of edge-cloud collaborative systems [liu2021mpcenabled]. Protocol optimization encompasses multiple layers with distinct considerations: (1) Application-layer protocols such as HTTP/HTTPS remain fundamental for web-based edge services, while MQTT and CoAP provide lightweight alternatives specifically designed for resource-constrained IoT devices with intermittent connectivity [phung2020enhancing]. MQTT’s publish-subscribe model offers excellent scalability for sensor data aggregation scenarios, whereas CoAP’s REST-based approach provides seamless integration with existing web infrastructures, though MQTT introduces broker dependency that can create single points of failure, while CoAP’s UDP-based nature may require additional mechanisms for reliable delivery in lossy networks [qu2024mobile]; (2) Network-layer protocols including IPv6 and emerging protocols like named data networking (NDN) address the scalability challenges of massive IoT deployments. IPv6 provides abundant address space essential for large-scale edge device connectivity, while NDN’s content-centric approach enables efficient data distribution and caching across edge networks, nevertheless IPv6 adoption remains inconsistent across heterogeneous edge environments, and NDN requires significant infrastructure changes that may limit near-term deployment [gundogan2020iot]; and (3) Transport-layer innovations such as QUIC and specialized protocols like BBR congestion control offer improved performance over traditional TCP in high-latency, variable-bandwidth edge-cloud connections [joarder2024exploring]. Specifically, QUIC’s built-in encryption and reduced connection establishment overhead particularly benefit mobile edge scenarios, whereas BBR’s model-based congestion control adapts better to the diverse network conditions encountered in edge deployments. Consequently, the choice of appropriate protocols depends on specific application requirements, device capabilities, and network characteristics, with optimal edge-cloud systems often employing protocol stacks that dynamically adapt to changing conditions and requirements [kawana2024communication].
The imperative for advanced model optimization and resource management in ECCC is driven by the demanding requirements of emerging application scenarios. For instance, autonomous driving requires ultra-low latency for real-time decision-making and must process vast streams of sensor data while preserving privacy [poularakis2019joint]. Augmented and virtual reality (AR/VR) applications demand high-bandwidth, low-latency rendering to deliver immersive user experiences, which is computationally intensive and energy-draining for mobile devices [vanhuynh2022edge]. Similarly, smart manufacturing relies on real-time fault diagnosis and predictive maintenance, necessitating reliable and continuous model adaptation in complex industrial environments [ren2021synergy]. Collectively, these applications create a tension between computational complexity, resource constraints, and performance requirements, motivating the development of the enabling technologies surveyed in this survey. A systematic mapping of these application scenarios to their technical challenges and enabling solutions is provided in the Supplementary Material (see LABEL:tab:app_summary in LABEL:sec:supp-application-scenarios), where we detail how different domains leverage the technologies surveyed in this paper.
@@ I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, each with distinct scopes and technical emphases. Early works like Abbas et al. [abbas2018mobile] and Ferrer et al. [ferrer2019decentralised] establish a foundational understanding of edge computing and decentralized cloud architectures respectively, with their primary focus on architectural foundations rather than AI-driven model optimization. Similarly, Liu et al. [liu2019survey] provide valuable insights into edge computing systems and tools, while Wang et al. [wang2019edge] contribute significantly to offloading algorithms, with their emphasis primarily on different aspects of edge-cloud integration rather than comprehensive AI technique integration. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [cao2021survey] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management, with their primary emphasis on system-level considerations rather than model optimization techniques. Domain-specific surveys like Arthurs et al. [arthurs2022taxonomy] focus on intelligent transportation systems with specialized depth in this domain. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain an active area of exploration, with Qu et al. [qu2025mobile] providing a dedicated investigation into mobile edge intelligence for LLMs, focusing primarily on resource-efficient techniques within their specialized scope. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [vanhuynh2022edge], offer valuable insights into ultra-reliable and low-latency communications with their emphasis on domain-specific considerations.
@@
The most relevant contemporary works include Yang et al. [yang2025edge], who extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, with their work emphasizing network dynamics as their primary contribution. Similarly, the comprehensive works by Yao et al. [yao2023edgecloud], Gu et al. [gu2024aienhanced], and Wang et al. [wang2024endedgecloud] each make valuable contributions with distinct emphases. Specifically, [yao2023edgecloud] provides excellent coverage of edge-cloud collaboration and AI techniques with particular strength in these core areas. [gu2024aienhanced] delivers strong analysis of AI-enhanced cloud-edge-terminal networks with good application coverage, emphasizing network architecture aspects. [wang2024endedgecloud] offers comprehensive model optimization coverage across the end-edge-cloud spectrum with particular depth in optimization techniques.
@@
Two surveys are particularly relevant to our work and warrant detailed comparison. Yang et al. [yang2025edge] provides a comprehensive survey of reinforcement learning approaches for mobile edge computing with excellent coverage of resource management and autonomous driving applications. Their work has a specific focus that differs from our survey in several key aspects: (1) Specialized computing paradigm emphasis, while they focus extensively on edge computing and resource management, our work provides comprehensive coverage of cloud integration and edge-cloud collaboration frameworks; (2) Domain-specific technical focus, their emphasis on reinforcement learning, while thorough, represents a specialized approach compared to our comprehensive treatment of model optimization techniques such as compression, adaptation, and neural architecture search that are central to modern ECCC systems; (3) Different scope for privacy and security, their work emphasizes other technical aspects, while our survey provides systematic attention to privacy-preserving mechanisms and security enhancement for real-world ECCC deployments; (4) Specialized application coverage, while strong in autonomous driving, our work provides comprehensive coverage of manufacturing and healthcare applications that increasingly rely on edge-cloud collaboration. Qu et al. [qu2025mobile] presents a focused investigation into mobile edge intelligence for large language models with strong coverage of LLMs, privacy considerations, and healthcare applications. Their work complements our survey with different emphases: (1) Specialized technical scope, while excellent in LLM-specific optimizations, our survey provides systematic coverage of general model optimization techniques applicable across diverse AI models and architectures; (2) Mobile edge intelligence focus, their focus on mobile edge intelligence represents a specialized area, while our work emphasizes comprehensive edge-cloud collaboration frameworks and distributed intelligence paradigms; (3) Specialized application domain, while strong in healthcare applications, our survey provides detailed analysis of manufacturing and industrial IoT scenarios that represent major application domains for ECCC; (4) Domain-specific security analysis, their privacy coverage addresses specific requirements, while our work provides comprehensive treatment of security mechanisms and adversarial considerations essential for robust ECCC systems.
@@
While the recent survey by [qu2025mobile] provides a valuable overview of AI models for edge computing, our work offers several distinct contributions. First, we present a more fine-grained taxonomy for model optimization techniques, particularly in the context of co-design with resource management (as detailed in Sec. III and Sec. IV). Second, our survey uniquely focuses on the interplay between task characteristics and adaptive AI strategies, a dimension not explicitly covered in [qu2025mobile]. Finally, we provide a more forward-looking perspective on open challenges, integrating insights from both system-level and algorithm-level research.
@@
[abbas2018mobile] 2018 Mobile Edge Comp., Network Architecture
[ferrer2019decentralised] 2019 Decentralized Cloud, Ad Hoc Computing
[liu2019survey] 2019 Edge Computing, Systems, Tools
[wang2019edge] 2019 Edge Cloud, Offloading Algorithms
[cao2021survey] 2021 Edge Computing, ECC, CPS
[arthurs2022taxonomy] 2022 EC Computing, ITS, Connected Vehicles
[vanhuynh2022edge] 2022 Edge Intel., URLLC, Digital Twin
[kar2023offloading] 2023 Offloading, Federated Cloud-Edge-Fog
[wang2023edge] 2023 Edge Computing, Sensor-Cloud
[yao2023edgecloud] 2023 EC Polarization, Collab. AI, Edge Comp.
[xu2024unleashing] 2024 AIGC, Mobile Networks, EC Generative AI
[gu2024aienhanced] 2024 AI, CET Network, Collaborative Computing
[wang2024endedgecloud] 2024 DL, Distributed Computing, Model Opt.
[yang2025edge] 2025 RL, Mobile Edge Comp., Res. Mgmt.
[qu2025mobile] 2025 Mobile Edge Intel., LLMs, Res. Mgmt.
@@ I-C Scope and Contributions
Figure 1: Overall structure of this survey, wherein the framework is organized into five layers: (1) Overview & architecture covering introduction and edge-cloud fundamentals, (2) technical implementation focusing on model optimization and resource management techniques, (3) protection framework addressing privacy and security enhancement, (4) practical deployment discussing applications and performance analysis, and (5) research vision outlining challenges and future directions.
As illustrated in Fig. 1, we provide a systematic and multi-layered examination of ECCC, addressing the critical intersection of distributed intelligence and model optimization. Edge-cloud systems face unprecedented challenges due to heterogeneous infrastructures, resource limitations, and growing computational demands from AI applications, making a comprehensive review of this rapidly evolving field both timely and necessary. To structure our comprehensive review, this survey is organized according to the five-layer framework depicted in Fig. 1, which reflects a top-down approach from foundational concepts to future research. Specifically, our organizational design provides a clear and logical progression through the complex ECCC landscape, starting with the Overview & Architecture layer that establishes the groundwork by covering introductory concepts and fundamental architectures. Building upon this foundation, the Technical Implementation layer delves into the core enabling technologies, including the model optimization and resource management strategies detailed in Sec. III and Sec. IV. Subsequently, the framework addresses security, practical deployment, and forward-looking research through dedicated layers, thereby ensuring a holistic examination of the field.
@@
By systematically analyzing ECCC architectures, resource management strategies, and performance metrics, the survey bridges existing gaps in understanding how AI can enhance decentralized computing environments. Moreover, it investigates critical enabling technologies such as model compression, adaptation, resource-aware optimization, and security mechanisms, which are pivotal for overcoming current limitations, spanning from architectural foundations and optimization techniques to practical applications and future research directions, thus offering researchers and practitioners a holistic perspective of this transformative paradigm. Regarding the three primary technical categories of ECCC, each addresses distinct challenges and exhibits specific applicability constraints. Model optimization techniques (Sec. III) primarily target the fundamental challenge of deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency. Model optimization approaches are particularly suitable for scenarios involving deep learning model deployment, privacy-sensitive distributed training, and real-time inference applications. However, their effectiveness is constrained by the inherent trade-offs between model accuracy and compression ratios, the heterogeneity of edge device capabilities, and the complexity of managing distributed learning processes. Although model compression methods excel in reducing computational overhead, they may compromise accuracy, whereas federated learning approaches provide privacy preservation at the cost of increased communication complexity and convergence challenges with non-IID settings.
AI-driven resource management strategies (Sec. IV) focus on optimizing resource utilization and task orchestration across the heterogeneous edge-cloud continuum, proving most effective in scenarios with dynamic workloads, multi-tenant environments, and applications requiring intelligent load balancing across diverse computational resources. Nevertheless, primary limitations include the computational complexity of multi-objective optimization problems, the difficulty in predicting system behavior under varying conditions, and scalability challenges in large-scale deployments. Although task offloading mechanisms can significantly reduce latency and energy consumption, they require sophisticated algorithms to handle dependency relationships and may introduce communication overhead that negates performance benefits. Furthermore, privacy and security enhancement frameworks (Sec. V) address the critical requirements for protecting sensitive data and ensuring system trustworthiness in distributed environments, particularly essential for healthcare, financial, and government applications where regulatory compliance and data protection are paramount. Nevertheless, privacy-preserving mechanisms often introduce computational overhead and complexity in key management, while security measures must continuously evolve to address emerging threats in heterogeneous edge environments. To provide a high-level overview of the core technical domains, Table II presents a comparative analysis of model optimization (Sec. III), AI-driven resource management (Sec. IV), and privacy and security enhancement (Sec. V). Specifically, this summary table evaluates each area across its core techniques, advantages, limitations, and applicable scenarios, offering a structured guide to the key trade-offs and design considerations discussed in the subsequent sections. The contributions of this survey are summarized as follows:
TABLE II: Comprehensive comparison of edge-cloud collaborative computing technical categories.
Category Core Techniques Advantages Limitations Applicable Scenarios
Model Optimization for EC Systems Advanced model compression techniques including pruning, quantization, and knowledge distillation, combined with adaptive learning approaches like transfer learning, federated learning, and continual learning for optimal deployment across resource-constrained edge-cloud infrastructures. Enables deployment of complex AI models on resource-constrained edge devices Reduces computational overhead and memory requirements significantly Maintains acceptable performance while optimizing for efficiency Supports adaptive learning and dynamic model selection Facilitates collaborative training without centralized data sharing Model compression may lead to accuracy degradation Federated learning faces challenges with non-IID data distributions Continual learning suffers from catastrophic forgetting issues NAS requires significant computational resources for architecture search Trade-offs between model accuracy, efficiency, and deployment constraints Deployment of deep learning models on IoT devices and edge nodes Privacy-sensitive applications requiring distributed training Dynamic environments with evolving data distributions Resource-constrained scenarios requiring efficient inference Real-time applications with strict latency requirements
AI-Driven Resource Management Intelligent task offloading mechanisms employing latency-aware, energy-efficient, and dependency-aware strategies, coupled with dynamic resource allocation and collaborative management approaches to optimize system performance and energy efficiency across distributed heterogeneous infrastructure. Optimizes resource utilization across heterogeneous infrastructure Enables intelligent task placement and load balancing Reduces overall system energy consumption and operational costs Supports dynamic adaptation to changing workload demands Facilitates collaborative resource sharing across edge-cloud continuum Complex optimization problems with multiple conflicting objectives Requires accurate prediction of system state and workload patterns Communication overhead between edge and cloud components Scalability challenges in large-scale distributed environments Difficulty in handling unpredictable workload fluctuations Large-scale IoT deployments with heterogeneous devices Time-critical applications requiring intelligent task scheduling Energy-constrained environments prioritizing sustainability Multi-tenant edge-cloud systems requiring fair resource sharing Dynamic workloads with varying computational demands
Privacy and Security Enhancement Comprehensive data privacy protection through privacy-preserving learning, secure multi-party computation, and differential privacy, integrated with model security mechanisms for integrity protection and attack detection, plus robust communication security using encryption and blockchain technologies. Protects sensitive data throughout the edge-cloud computing pipeline Enables secure collaborative learning and inference Provides robust defense mechanisms against various attack vectors Ensures compliance with privacy regulations and standards Maintains data integrity and system trustworthiness Privacy-preserving techniques often introduce computational overhead Trade-offs between security level and system performance Complex key management and authentication in distributed systems Potential vulnerabilities in heterogeneous edge devices Evolving threat landscape requiring continuous security updates Healthcare and financial applications with strict privacy requirements Multi-party collaborative learning scenarios Industrial IoT systems handling sensitive operational data Government and defense applications requiring high security Cross-domain data sharing with regulatory compliance needs
@@
  • First Comprehensive Survey. To the best of our knowledge, this is the first comprehensive survey that specifically targets the intersection of distributed intelligence and model optimization within the broader field of edge-cloud collaborative computing, reviewing recent advances across architectures, techniques, security, and privacy.